OpenStudy (anonymous):

Jackson is playing a video game that has 4 castles to conquer within each level. Jackson has completed 6 levels so far. If x represents the number of remaining levels to complete, which of the following equations can be used to find the total number of castles Jackson will have conquered once he completes the final level

OpenStudy (anonymous):

wouldnt it be \[y=6x+4\]

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Koikkara help plz

OpenStudy (anonymous):

or is it \[y=24x+4\]

OpenStudy (michele_laino):

the castles conquered are 4*6=24, so the requested function is: \[\Large y = 24 + 4x\] where x is the number of remaining levels

OpenStudy (anonymous):

there not an equation that looks like that

OpenStudy (anonymous):

the only one that looks like that is the one i posted

OpenStudy (anonymous):

@nomymalik

OpenStudy (michele_laino):

has Jason completed 6 levels?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

so, he has conqured 6*4= 24 castels, right?

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

LET ME WORK IT OUT

OpenStudy (anonymous):

LMAO OK NOMY

OpenStudy (michele_laino):

how many levels are in total?

OpenStudy (anonymous):

it doesnt say

OpenStudy (anonymous):

go to math ways its a math website

OpenStudy (anonymous):

send me a link plz

OpenStudy (anonymous):

https://mathway.com/

OpenStudy (anonymous):

@nomymalik what do i do now

OpenStudy (anonymous):

put the problem

OpenStudy (anonymous):

th equation

OpenStudy (anonymous):

@nomymalik

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@nomymalik that didnt help

OpenStudy (koikkara):

@reneesweet17 QH can Help you learn things in a correct way, He is an experienced user. Listen to Him !

OpenStudy (michele_laino):

I think that the right equation is: \[y = 24 + 4x\]

OpenStudy (anonymous):

oh then idk

OpenStudy (michele_laino):

as I explained before

OpenStudy (anonymous):

likie i said the only one that look even close to that one is the one i posted on the comments

OpenStudy (michele_laino):

I think that if there are 4 catles to conquer in a single level, and Jason has completed 6 levels, then Jason has conquered 6*4 = 24 castels. Now if Jason has to complete x levels, then when he has completed those x levels, he has conquered: 24+ 4*x castles

OpenStudy (koikkara):

\(Sample~Question\) . Jackson is playing a video game that has 3 castles to conquer within each level. Jackson has completed 5 levels so far. If x represents the number of remaining levels to complete, which of the following equations can be used to find the total number of castles Jackson will have conquered once he completes the final level? \(Solution\) Jackson has already completed 5 levels with 3 castles conquered per level. So, he has already conquered 15 castles. Each remaining level, represented by x, contains 3 castles to conquer. Therefore, the equation that can be used to find the total number of castles Jackson will have conquered once he completes the final level is shown below. y = 3x + 15 @reneesweet17 Does this look similar to your question ? \(try..how~y=3x~+~15\) came... and you can sort out your question !! \(Hope~it~helps~!\)

OpenStudy (anonymous):

im not even on the same one it was the one i posted